OpenStudy (anonymous):

Directions: Click on "2nd Level". How can you prove that b - a is the negative of a - b?

OpenStudy (anonymous):

do u kno geometry

OpenStudy (anonymous):

no this is in my common core algebra class

OpenStudy (amistre64):

common core eh ... you cant prove nothing in common core :/

OpenStudy (anonymous):

its wat my class is called

OpenStudy (amistre64):

-(a-b) = b-a 0 = (b-a) + (a-b) 0 = b +(-a + a) -b 0 = b +0 -b 0 = b -b 0 = 0, is ture

OpenStudy (amistre64):

of course you would have to name each step according to its mathical property

OpenStudy (anonymous):

what is (2xy) times (9x+9)

OpenStudy (anonymous):

i think it is 18xy+18xy but im not sure